OpenStudy (anonymous):

Find the slope and the y-intercept of the equation x − 4(y + 1) − 5 = 0.

OpenStudy (anonymous):

First, put it into slope intercept form

OpenStudy (anonymous):

I don't know how to that

OpenStudy (anonymous):

First you simplify it so... x - 4y - 4 - 5=0 So using the distributive property.

OpenStudy (anonymous):

whats the next step to do?

OpenStudy (anonymous):

Simplify the numbers so do -4 - 5 and write it out.

OpenStudy (anonymous):

i got -9 for that

OpenStudy (anonymous):

Good!

OpenStudy (anonymous):

so whats next

OpenStudy (anonymous):

So, now you have x - 4y - 9=0. We have to get the y alone. So, now we add x to both sides.

OpenStudy (anonymous):

After, we have -4y - 9= -x

OpenStudy (anonymous):

thats what i got lol

OpenStudy (anonymous):

Good!

OpenStudy (anonymous):

Now, we need to add 9 to both sides. To get 4y = -x + 9

OpenStudy (anonymous):

is there still a negative in front of the 4y?

OpenStudy (anonymous):

Yes, sorry

OpenStudy (anonymous):

We have now -4y = -x + 9

OpenStudy (anonymous):

okay. lol

OpenStudy (anonymous):

i got -13

OpenStudy (anonymous):

Next, we divide everything by 4.

OpenStudy (anonymous):

i think im doing it wrong..

OpenStudy (anonymous):

It is okay.

OpenStudy (anonymous):

so whats the next step?